# OccFusion: A Straightforward and Effective Multi-Sensor Fusion Framework for 3D Occupancy Prediction

### 03 Mar 2024 v1

@@ IV-D Model Performance Analysis
Method Backbone Input Modality Resolution mIoU $\bullet$ barrier $\bullet$ bicycle $\bullet$ bus $\bullet$ car $\bullet$ const. veh. $\bullet$ motorcycle $\bullet$ pedestrian $\bullet$ traffic cone $\bullet$ trailer $\bullet$ truck $\bullet$ drive. surf. $\bullet$ other flat $\bullet$ sidewalk $\bullet$ terrain $\bullet$ manmade $\bullet$ vegetation MonoScene [1] R101-DCN C $200\times 200\times 16$ 7.31 4.03 0.35 8.00 8.04 2.90 0.28 1.16 0.67 4.01 4.35 27.72 5.20 15.13 11.29 9.03 14.86 Atlas [62] R101-DCN C $200\times 200\times 16$ 15.00 10.64 5.68 19.66 24.94 8.90 8.84 6.47 3.28 10.42 16.21 34.86 15.46 21.89 20.95 11.21 20.54 BEVFormer [3] R101-DCN C $200\times 200$ 16.75 14.22 6.58 23.46 28.28 8.66 10.77 6.64 4.05 11.20 17.78 37.28 18.00 22.88 22.17 13.80 22.21 TPVFormer [5] R101-DCN C $200\times 200\times 16$ 17.10 15.96 5.31 23.86 27.32 9.79 8.74 7.09 5.20 10.97 19.22 38.87 21.25 24.26 23.15 11.73 20.81 SurroundOcc [7] R101-DCN C $200\times 200\times 16$ 20.30 20.59 11.68 28.06 30.86 10.70 15.14 14.09 12.06 14.38 22.26 37.29 23.70 24.49 22.77 14.89 21.86 InverseMatrixVT3D [14] R101-DCN C $200\times 200\times 16$ 25.95 26.13 15.19 42.65 43.92 11.42 20.82 22.68 13.45 22.33 36.04 47.47 16.26 21.47 25.33 16.37 33.73 OccFusion R101+VoxelNet C+R $200\times 200\times 16$ 28.07 27.82 17.71 47.41 46.60 13.73 23.02 22.72 17.77 26.29 37.58 47.56 19.45 27.03 23.88 22.90 27.56 OccFusion R101+VoxelNet C+L $200\times 200\times 16$ 33.92 31.79 18.91 49.36 48.95 24.77 24.10 29.70 22.26 29.91 36.09 47.98 24.18 28.18 31.87 40.82 53.94 OccFusion R101+VoxelNet C+L+R $200\times 200\times 16$ 34.77 33.05 22.46 46.62 52.51 22.73 26.96 30.42 20.74 25.42 41.24 53.47 19.83 31.51 35.56 39.25 54.57
@@ IV-E Challenging Scenarios Performance Analysis
Method Backbone Input Modality Resolution mIoU $\bullet$ barrier $\bullet$ bicycle $\bullet$ bus $\bullet$ car $\bullet$ const. veh. $\bullet$ motorcycle $\bullet$ pedestrian $\bullet$ traffic cone $\bullet$ trailer $\bullet$ truck $\bullet$ drive. surf. $\bullet$ other flat $\bullet$ sidewalk $\bullet$ terrain $\bullet$ manmade $\bullet$ vegetation SurroundOcc [7] R101-DCN C $200\times 200\times 16$ 19.85 21.40 12.75 25.49 31.31 11.39 12.65 8.94 9.48 14.51 21.52 35.34 25.32 29.89 18.37 14.44 24.78 InverseMatrixVT3D [14] R101-DCN C $200\times 200\times 16$ 27.94 25.08 14.30 43.25 43.32 14.46 21.29 15.93 12.48 28.82 36.40 55.31 23.66 37.75 22.53 18.38 34.06 OccFusion R101+VoxelNet C+R $200\times 200\times 16$ 28.19 27.59 18.73 43.24 48.67 12.05 18.42 15.85 14.46 29.68 38.03 53.62 19.05 35.38 20.89 23.86 31.47 OccFusion R101+VoxelNet C+L $200\times 200\times 16$ 33.24 27.77 17.91 46.36 48.90 22.57 23.43 23.53 22.54 33.75 36.26 49.09 25.53 34.70 23.41 40.58 55.58 OccFusion R101+VoxelNet C+L+R $200\times 200\times 16$ 34.35 29.06 21.53 44.83 54.01 20.00 23.41 23.82 22.42 28.85 42.22 58.41 20.86 38.55 24.57 39.96 57.04
@@
Method Backbone Input Modality Resolution mIoU $\bullet$ barrier $\bullet$ bicycle $\bullet$ bus $\bullet$ car $\bullet$ const. veh. $\bullet$ motorcycle $\bullet$ pedestrian $\bullet$ traffic cone $\bullet$ trailer $\bullet$ truck $\bullet$ drive. surf. $\bullet$ other flat $\bullet$ sidewalk $\bullet$ terrain $\bullet$ manmade $\bullet$ vegetation SurroundOcc [7] R101-DCN C $200\times 200\times 16$ 10.80 10.55 14.60 0.00 31.05 0.00 8.26 5.37 0.58 0.00 18.75 30.72 2.74 12.39 11.53 10.52 15.77 InverseMatrixVT3D [14] R101-DCN C $200\times 200\times 16$ 12.68 16.33 6.68 0.00 38.88 0.00 15.36 8.07 1.06 0.00 24.24 33.32 2.64 11.64 10.30 4.28 17.37 OccFusion R101+VoxelNet C+R $200\times 200\times 16$ 15.39 12.93 12.50 0.00 48.28 0.00 18.74 9.12 1.28 0.00 29.39 41.57 2.01 18.53 11.63 13.59 26.73 OccFusion R101+VoxelNet C+L $200\times 200\times 16$ 20.77 14.24 13.84 0.00 47.04 0.00 17.67 13.25 1.17 0.00 27.26 44.47 2.54 17.89 28.72 32.36 51.10 OccFusion R101+VoxelNet C+L+R $200\times 200\times 16$ 21.41 17.36 10.66 0.00 51.52 0.00 18.47 14.84 0.68 0.00 31.52 42.83 2.10 18.17 30.86 29.40 52.69

### 13 Mar 2024 v2

@@ IV-D Model Performance Analysis
Method Backbone Input Modality Resolution mIoU $\bullet$ barrier $\bullet$ bicycle $\bullet$ bus $\bullet$ car $\bullet$ const. veh. $\bullet$ motorcycle $\bullet$ pedestrian $\bullet$ traffic cone $\bullet$ trailer $\bullet$ truck $\bullet$ drive. surf. $\bullet$ other flat $\bullet$ sidewalk $\bullet$ terrain $\bullet$ manmade $\bullet$ vegetation InverseMatrixVT3D [14] R101-DCN C $200\times 200\times 16$ 25.95 26.13 15.19 42.65 43.92 11.42 20.82 22.68 13.45 22.33 36.04 47.47 16.26 21.47 25.33 16.37 33.73 OccFusion R101+VoxelNet C+R $200\times 200\times 16$ 28.07 27.82 17.71 47.41 46.60 13.73 23.02 22.72 17.77 26.29 37.58 47.56 19.45 27.03 23.88 22.90 27.56 OccFusion R101+VoxelNet C+L $200\times 200\times 16$ 33.92 31.79 18.91 49.36 48.95 24.77 24.10 29.70 22.26 29.91 36.09 47.98 24.18 28.18 31.87 40.82 53.94 OccFusion R101+VoxelNet C+L+R $200\times 200\times 16$ 34.77 33.05 22.46 46.62 52.51 22.73 26.96 30.42 20.74 25.42 41.24 53.47 19.83 31.51 35.56 39.25 54.57
@@ IV-E Challenging Scenarios Performance Analysis
Method Backbone Input Modality Resolution mIoU $\bullet$ barrier $\bullet$ bicycle $\bullet$ bus $\bullet$ car $\bullet$ const. veh. $\bullet$ motorcycle $\bullet$ pedestrian $\bullet$ traffic cone $\bullet$ trailer $\bullet$ truck $\bullet$ drive. surf. $\bullet$ other flat $\bullet$ sidewalk $\bullet$ terrain $\bullet$ manmade $\bullet$ vegetation InverseMatrixVT3D [14] R101-DCN C $200\times 200\times 16$ 27.94 25.08 14.30 43.25 43.32 14.46 21.29 15.93 12.48 28.82 36.40 55.31 23.66 37.75 22.53 18.38 34.06 OccFusion R101+VoxelNet C+R $200\times 200\times 16$ 28.19 27.59 18.73 43.24 48.67 12.05 18.42 15.85 14.46 29.68 38.03 53.62 19.05 35.38 20.89 23.86 31.47 OccFusion R101+VoxelNet C+L $200\times 200\times 16$ 33.24 27.77 17.91 46.36 48.90 22.57 23.43 23.53 22.54 33.75 36.26 49.09 25.53 34.70 23.41 40.58 55.58 OccFusion R101+VoxelNet C+L+R $200\times 200\times 16$ 34.35 29.06 21.53 44.83 54.01 20.00 23.41 23.82 22.42 28.85 42.22 58.41 20.86 38.55 24.57 39.96 57.04
@@
Method Backbone Input Modality Resolution mIoU $\bullet$ barrier $\bullet$ bicycle $\bullet$ bus $\bullet$ car $\bullet$ const. veh. $\bullet$ motorcycle $\bullet$ pedestrian $\bullet$ traffic cone $\bullet$ trailer $\bullet$ truck $\bullet$ drive. surf. $\bullet$ other flat $\bullet$ sidewalk $\bullet$ terrain $\bullet$ manmade $\bullet$ vegetation InverseMatrixVT3D [14] R101-DCN C $200\times 200\times 16$ 12.68 16.33 6.68 0.00 38.88 0.00 15.36 8.07 1.06 0.00 24.24 33.32 2.64 11.64 10.30 4.28 17.37 OccFusion R101+VoxelNet C+R $200\times 200\times 16$ 15.39 12.93 12.50 0.00 48.28 0.00 18.74 9.12 1.28 0.00 29.39 41.57 2.01 18.53 11.63 13.59 26.73 OccFusion R101+VoxelNet C+L $200\times 200\times 16$ 20.77 14.24 13.84 0.00 47.04 0.00 17.67 13.25 1.17 0.00 27.26 44.47 2.54 17.89 28.72 32.36 51.10 OccFusion R101+VoxelNet C+L+R $200\times 200\times 16$ 21.41 17.36 10.66 0.00 51.52 0.00 18.47 14.84 0.68 0.00 31.52 42.83 2.10 18.17 30.86 29.40 52.69